1. Shea is walking on a circular path. The path has a diameter of 8 miles. If

Shea walks about 4 miles an hour, how many hours will it take him to

complete the path? Round your answer to the nearest hour.





1
Expert's answer
2021-01-20T17:26:32-0500

Here Shea is walking on a circular path of diameter 8miles.


So, radius of the path is "r=4miles"


So, circumference of the circular path is:-



"C=2\pi r\\
\Rightarrow C=2\cdot3.14\cdot4miles=25.12miles"

So, total distance to cover is 25.12miles.


Also we know that Shea walks at around 4miles/hr.


So, total time taken to complete the path:-



"t=\frac{C}{v}\\~\\
\Rightarrow t=\frac{25.12}{4}\\~\\
\Rightarrow t=6.28\approx6.3hrs\approx6hrs"

So, time taken by Shea to walk is 6hrs.